OpenStudy (anonymous):

Vector question: a=2i+6j+10k b=4i+7j=2k Find two vectors u and v such that: v is perpendicular to a u is parallel a b=u+v

OpenStudy (experimentx):

v perpendicular to a, v.a = 0 u parallel to b , u = k(b) ... you can easily construct a vector ... is that last equation is also a requirement?

OpenStudy (experimentx):

sorry u = k(a)

OpenStudy (experimentx):

let

OpenStudy (experimentx):

v = xi+yj+zk, then 2x+6y+10z = 0, ----1 4i+7j-2k = xi+yj+zk + k (2i+6j+10k) --- 2 compare the coefficients on 2, put the value of x,y,z on 1 (in terms of k), find the value of k, and back to 2, find x, y, z

OpenStudy (experimentx):

i leave the rest of work to you

OpenStudy (anonymous):

ok thanks for your help